OpenStudy (anonymous):

(7,0) and (0-4) find the slope. Can some please explain this to me

OpenStudy (imstuck):

The slope formula is\[\frac{ y _{2} -y _{1}}{x _{2}-x _{1} }\]

OpenStudy (imstuck):

The means you subtract the y in your first point, which is 0, from the y in your second point, which is -4, and divide it by the x in your first point, which is 7, from the x in your second point, which is 0, and then simplify the result. In practice, it will look like this:

OpenStudy (imstuck):

\[\frac{ -4-0 }{ 0-7 }=\]

OpenStudy (imstuck):

That equals \[\frac{ -4 }{ -7 }=\frac{ 4 }{ 7 }\]In essence this means that the steepness of the line is positive, meaning it goes from the lower left in your plane to the upper right, and it does this at a steady rate of up 4 units and over 7 units to the right.

OpenStudy (anonymous):

Thank you so much I am really struggle with this and it needed to understand in order to get it....